OpenStudy (anonymous):

Use De Moivre's Theorem to find all of the complex solutions of the equation x³-i=0. Express the answer in rectangular form. Any help at all would be greatly appreciated. Thanks!

OpenStudy (anonymous):

x^3 = i = e^(i(pi/2 + 2kpi)) x = e^(i(pi/2 + 2KPI)/3)

OpenStudy (anonymous):

so x = e^(ipi/6), e^(5ipi/6) and e^(7ipi/6)

OpenStudy (anonymous):

so now expand these into rectangular form

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no, thank you!